The magnetic field of an electromagnetic wave in a vacuum is Bz=(3.0μT)sin((1.4×107)x−2πft), where x is in m and t is in s.

What is the wave's wavelength?

What is the wave's frequency?

What is the wave's electric field amplitude?
Answer #1

Solution)

Given Bmax=3.0 uT

Wave constant, k=1.4*10^7 m^-1

Part a) Wavelength, lambda= 2pi/k=2*3.14/(1.4*10^7)=4.48*10^-7 m

=======

Part b)

Frequency, f= c/ lambda= (3*10^8)/(4.48*10^-7)= 6.68*10^14 Hz

======

Part c)

Emax= cBmax

=(3*10^8)*(3*10^-6)= 900 N/C

=========